OpenStudy (anonymous):

Which point is a solution of x + 2y ≤ 4? (2, 4) (1, 1) (3, 5) (-1, 5)

OpenStudy (anonymous):

Just substitute values for x and y! :)

OpenStudy (anonymous):

try every point!

OpenStudy (anonymous):

\[(x,y)=(2,4)=2+2(4)\le4?\]

OpenStudy (anonymous):

is \[10\le4?\] clearly no :) that' how you do it

OpenStudy (anonymous):

I cant do it that's why I asked for help

OpenStudy (anonymous):



OpenStudy (anonymous):

(1,1) so x = 1, y =1 --- plugging it in to our equation we get (1) + 2(1) <= 4?

OpenStudy (anonymous):

(3,5) so x =3 , y =5 --- plugging it in to our equation we get (3) + 2(5) <= 4?

OpenStudy (anonymous):

(-1,5) so x = -1, y =5 --- plugging it in to our equation we get (-1) + 2(5) <= 4?